OpenStudy (anonymous):

Find the inverse of f(x) and name it g(x) f(x) = 1/3x+2

OpenStudy (freckles):

I assume you mean f(x)=1/(3x+2) or is it f(x)=1/3x+2?

OpenStudy (aum):

\[ f(x) = \frac 13x + 2 ? \\ f(x) = \frac{1}{3x} + 2 ? \\ f(x) = \frac{1}{3x+2} ? \]

OpenStudy (freckles):

@jwilliams03 which ever one that @aum posted all of those are actually one to one... So they each have an inverse function. You find the inverse function by solving for x. Then replace y with x and x with g(x) after the completion of the solving for x thing.

OpenStudy (aum):

jwilliams: My earlier reply is not the solution. I am asking for clarification. It is not clear which one of those three is the function that you are asked to find the inverse of. Can you answer the question?

OpenStudy (anonymous):

@aum it is the first one. I think i got the answer though. g(x) = 3x-6?

OpenStudy (aum):

correct. g(x) = 3x - 6